OpenStudy (anonymous):

Find the first derivative of y=(x+x^-1)^2

OpenStudy (perl):

did you try chain rule

OpenStudy (perl):

y = u^n dy/dx = n * u^(n-1) * du/dx

OpenStudy (anonymous):

Yeah i tried it but i keep getting an answer that is not on the matching answer sheet so i think im doing something wrong.

OpenStudy (perl):

dy/dx = 2 * ( x + x^-1 )^1 * ( 1 - x^-2 )

OpenStudy (jhannybean):

\[f(x) = (x+x^{-1})^2 = \left(x+\frac{1}{x}\right)^2 = \left(\frac{x^2+1}{x}\right)^2\]

OpenStudy (jhannybean):

That might make things easier, maybe.

OpenStudy (anonymous):

Here is how I like to look at it: \[ \begin{split} dy &= d[(x+x^{-1})^2] &\text{let}\;u=(x+x^{-1}) \\ &=d(u^2)\\ &= 2udu\\ &= 2(x+x^{-1})d(x+x^{-1}) \end{split} \]

OpenStudy (anonymous):

It has to be simplified farther than that and the chain rule has to be used but thank you! I keep getting confused when i try to simplify it past what Perl got.

OpenStudy (jhannybean):

\[f'(x) = 2\left(\frac{x^2+1}{x}\right) \cdot \left[\frac{2x(x)-1(x^2+1)}{x^2}\right]\]\[f'(x) =2\left(\frac{x^2+1}{x}\right)\cdot\left[\frac{2x^2-x^2-1}{x^2}\right]\]

OpenStudy (jhannybean):

That way work.

OpenStudy (jhannybean):

Then just simplify what is in the brackets (my explanation) Albeit, there are various methods of solving something as trivial as this question.

OpenStudy (anonymous):

Here is my complete version: \[ \begin{split} dy &= d[(x+x^{-1})^2] &\text{let}\;u=(x+x^{-1}) \\ &=d(u^2)\\ &= 2udu\\ &= 2(x+x^{-1})d(x+x^{-1})\\ &=2(x+x^{-1})(1-x^{-2})dx \end{split}\\ \; \\ \implies \frac{dy}{dx} = 2(x+x^{-1})(1-x^{-2}) \]

OpenStudy (jhannybean):

If you simplify @wio 's answer in making all variables positive, you will end up with my answer, after you have simplified it. :-) I love derivatives.